OpenStudy (anonymous):

HELP PLEASE ? Checking account A charges a monthly service fee of $12 and a wire transfer fee of $10.50, while checking account B charges a monthly service fee of $21 and a wire transfer fee of $8.50. Which checking account is the better deal if four wire transfers are made per month?

undeadknight26 (undeadknight26):

F(x) = 12x + 10.50 F(x) = 21x + 8.60

undeadknight26 (undeadknight26):

now plug in 4 and solve

undeadknight26 (undeadknight26):

wait no i did the equations wrong. f(x) = 10.50x + 12 F(x) = 8.50 + 21

undeadknight26 (undeadknight26):

so now the first one would be F(4) = 10.50 x 4 + 12 f(4) = 42 + 12 f(4) = 54